after a typhoon, the lineman needed 1500 meters of electric wire for the restoration of electricity in a village. He divided the wire into 5 pieces in the ratio 1:2:3:4:5. How long is each piece of electric wire?

Answer by ikleyn(52780)

.

The pieces are  x, 2x, 3x, 4x and 5x meters long,  where "x" is the common measure, now unknown.


To find x, use the equation for the total length


    x + 2x + 3x + 4x + 5x = 1500 meters.


Simplify and find x


    15x = 1500

      x = 1500/12 = 100.


ANSWER.  The pieces are  100 m, 200 m, 300 m, 400 m and 500 m long.

Solved.